OpenStudy (anonymous):

Topology in R^n! Given J={(x,y):x^-y<=1}, the int J={x^2-y<1} and the J'=[x^2-y<=1}. It's not-open, closed, and not-limited! Can someone explain me why it is not limited, if it can't be more than 1?

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

so u understand why its not open and closed ?

OpenStudy (loser66):

I know why it is not open but don't know why it is not closed. Please, explain

OpenStudy (anonymous):

its closed :O

OpenStudy (loser66):

oh, it is not-open, closed. Mean it is not-open and it is closed. :)

OpenStudy (loser66):

ok, makes sense. How about not-limited?

OpenStudy (anonymous):

j is closed since j' in topology

OpenStudy (loser66):

Not that, right? it is closed because \(J^c\) is open

OpenStudy (anonymous):

well not really , sometimes a set is open and closed at the same time , also it might be not closed not open ( according to the definition ) yes exactly lol but see the question j' is j^c :)

OpenStudy (anonymous):

and open means j in topology

OpenStudy (loser66):

How? in J', we have equal sign there --> it is not \(J^c\)

OpenStudy (loser66):

It is \(J^c \cup bd J\)

OpenStudy (loser66):

= Ext (J) , right?

OpenStudy (anonymous):

ohh :O yes right , sorry :P ill the question dint specify things

OpenStudy (loser66):

However, I don't understand why int J = {x^2 -y <1} while J ={x^-1 y \(\leq 1\)}

OpenStudy (anonymous):

I'm sorry for the delay on the reply. I was lunching. the int of J is the set of point that belongs only to the interior of J. or, as my professor explained it's any element of J that, a "circle" centered on it,should only contain elements of J

OpenStudy (anonymous):

I don't understand why J is not limited if it can't assume number less than 1.

OpenStudy (anonymous):

whats the definition of limited and not limited /?

OpenStudy (anonymous):

It's the existence of an upper or lower bond of the function?!?!

OpenStudy (anonymous):

are u sure ? im just asking wanna know if its the same in topology or not xD

OpenStudy (anonymous):

not sure what you're going with this..... please explain me like I never heard it before! ;)

OpenStudy (anonymous):



OpenStudy (anonymous):

hehe , im not sure of the definition of not limited in topology thats all :)

OpenStudy (anonymous):

but, by your example, the function is limited. couldn't take values minor than the function. \[x^2 - y \le 1\] Thus, is limited. It as a upper bound, right?

OpenStudy (anonymous):

if we are talking about functions then yes its limited , but idk if its the same in topology

OpenStudy (anonymous):

yes, we're talking about topology of functions in multivariable calculus. This exercise is about the interior, the interior+border, open, closed, limited, etc. of subsets of \[R ^{2}\]